OpenStudy (calculusxy):

Medal for the best answer and explanation. D) Find a number that has exactly 1. two factors 2. three factors 3. four factors 4. six factors 5. What is it special about these numbers? Explain.

OpenStudy (calculusxy):

Answers for questions 1-4 (I write) 1. 2 2. 4 3. 6 4. 12

ganeshie8 (ganeshie8):

oh u counted 1 also as a factor...

ganeshie8 (ganeshie8):

and the number itself, ok... looks legit :)

OpenStudy (calculusxy):

So how can I find the specialty of those numbers?

ganeshie8 (ganeshie8):

they're sum of previous numbers, is that some clue hmm

ganeshie8 (ganeshie8):

1. 2 2. 4 3. 6 = 4+2 4. 12 = 6+4+2

OpenStudy (calculusxy):

Thank you very much!

ganeshie8 (ganeshie8):

np :) thats all the q about ?

OpenStudy (calculusxy):

I can have more, but for now I don't have any.

ganeshie8 (ganeshie8):

also, are these the oly numbers that can have factors >= n/2 for ex, 12 has 6 factors : 12/2 = 6 6 has 4 factors : 6/2 = 3 ...

ganeshie8 (ganeshie8):

gimme few minutes. il confirm, if we cannot have any other +ve integers that can have >= n/2 factors

ganeshie8 (ganeshie8):

1. 2 2. 4 3. 6, 8 4. 12

ganeshie8 (ganeshie8):

8 also has 4 factors ! so speciality about these numbers is this :- Only these numbers can have factors >= n/2. Its not possible for other numbers